OpenStudy (anonymous):

Factor the Trinomial 8x2 − 10xy − 3y2 I can't seem to get the correct answer to this.

OpenStudy (mertsj):

8(-3)=-24 -12(2)=-24 AND -12+2=-10 Replace -10xy with -12xy+2xy and factor by grouping

OpenStudy (rushwr):

\[8x ^{2} -10xy -3y ^{2}\] \[8x ^{2} -12xy-2xy-3y ^{2}\] \[4x(2x-3y) +y(2x-3y)\] \[(4x+y) (2x-3y)\]